# Position: Meaning Is Not A Metric: Using LLMs to make cultural context legible at scale

### v1

@@ Abstract
This position paper argues that large language models (LLMs) can make cultural context, and therefore human meaning, legible at an unprecedented scale in AI-based sociotechnical systems. We argue that such systems have previously been unable to represent human meaning because they rely on thin descriptions: numerical representations that enforce standardization and therefore strip human activity of the cultural context that gives it meaning. By contrast, scholars in the humanities and qualitative social sciences have developed frameworks for representing meaning through thick description: verbal representations that accommodate heterogeneity and retain contextual information needed to represent human meaning. While these methods can effectively codify meaning, they are difficult to deploy at scale. However, the verbal capabilities of LLMs now provide a means of (at least partially) automating the generation and processing of thick descriptions, potentially overcoming this bottleneck. We argue that the problem of rendering human meaning legible is not just about selecting better metrics, but about developing new representational formats (based on thick description). We frame this as a crucial direction for the application of generative AI and identify five key challenges: preserving context, maintaining interpretive pluralism, integrating perspectives based on lived experience and critical distance, distinguishing qualitative content from quantitative magnitude, and acknowledging meaning as dynamic rather than static. Furthermore, we suggest that thick description has the potential to serve as a unifying framework to address a number of emerging concerns about the difficulties of representing culture in (or using) LLMs.
@@ 1 Introduction
The premise of this paper is that it is important, but difficult, to develop technologies that can support meaningful human experience. Our position aims to articulate how AI technologies can be used not just to yield superficial increases in convenience, but to offer something deeper: a tool for helping people tap more directly into what really matters. This is not an inevitable outcome of technological progress. Dystopian fiction is full of cases in which a hypothetical society has advanced far beyond our own, only to find that their sophisticated technology has had a detrimental effect on human experience. Of course, the easy part is proposing that technology should have positive societal impacts; the difficult part is defining “what really matters.”
For example, consider the case of social media. Initially these platforms were presented as a tool for human social connection [10]. Over time, however, they came to reflect the more superficial dynamics of engagement metrics; content was prioritized not to facilitate more profound social interactions, but on the basis of likes, views, and shares [103, 30, 63]. This led to a wide range of negative societal outcomes, such as rampant misinformation [55, 106], increased feelings of loneliness [78, 101], addictive behavior [40], and other psychological and social costs [22].
In response, technologists, social scientists, and activists have proposed other ways to prioritize content on these platforms which would, in one form another, allow us to more directly target what really matters [34, 97, 19]. A representative example is the Time Well Spent movement (now the Center for Humane Technology) led by Tristan Harris [34]. The goal of this movement was to develop technology that supports “meaningful” social connections and “choices we cherish” [72]. In response, at least one major social media platform alleged to have changed their recommendation algorithm to prioritize these kind of “meaningful” interactions [96]. However, it was disputed whether this effort succeeded, or simply substituted one superficial metric for another [72, 63].
@@
We contend that in order to develop sociotechnical systems that can increase time well spent or reduce drudgery, we first need a way of systematically codifying what kind of experiences humans find meaningful. We approach this problem—how to design sociotechnical systems that support human meaning—as one of representation. The core of our argument is that this issue has been difficult to address not simply because we have been using the wrong metrics, but because “metrics” (as we have traditionally conceived of them) are the wrong representational format. The formats we typically use for measuring behavior at scale exclude key contextual information necessary for representing human meaning. The primary aim of this paper will be to describe why that is the case and how we can begin to address it.
Our position is that we can use LLMs to make human meaning legible at an unprecedented scale; that “thin” metrics alone will not be able to do this, because they cannot adequately represent cultural context; and that developing “thick” representational formats might eventually allow our society’s most powerful and impactful sociotechnical systems to support the heterogeneous range of experiences that really matter to people. We argue that such systems have previously been unable to represent human meaning because they typically rely on thin descriptions: numerical representations that enforce standardization, and therefore strip human activity of the broader context that gives it meaning [9, 71, 59, 89]). By contrast, scholars in the humanities and qualitative social sciences have developed robust frameworks for representing meaning based on thick descriptions: verbal representations that allow for heterogeneity, and therefore retain crucial contextual information necessary for analyzing cultural and experiential meaning [27, 108, 87, 82]. These methods provide a way of codifying meaning—but they are difficult to deploy at scale.
@@ 2 The Legibility of Meaning
In his classic political science text Seeing like a State, James C. Scott (1998) introduces the concept of legibility [89]. This term describes the connection between a sociotechnical system (in Scott’s case, a political state) and its information processing capacities. Political states cannot represent the goals and desires of their constituents directly. Simply asking each person what they want would result in too much heterogeneity of response. Rather, the state must develop standardized metrics that can be applied across the whole population. These help the state understand the overall trends and dynamics of the populace. However, as Scott argues, by enforcing this standardization the state makes some kinds of information legible—while rendering other kinds illegible.
Drawing on Scott’s concept, we frame the main challenge of this paper as how to make human meaning legible at scale. To clarify: we are not concerned with how to make human meaning legible within AI models, but rather how AI models can be used to render new kinds of information legible within sociotechnical systems. We use the term sociotechnical system to refer to any large-scale configuration of people and machines that cannot be understood in terms of human or technological dynamics alone but that depend on a complex entanglement of their interactions [41, 5, 100]. This includes many systems that have a major impact on society, such as political states, social media platforms, corporations, and institutions for healthcare, education, and science. As AI technologies become increasingly integrated into these societally impactful systems, how can we use these technologies to make new kinds of information (such as human meaning) legible at scale?
@@ 2.1 What Do We Mean By “Meaning”?
The term “meaning” has been used by many different people, across many disciplines, to refer to many different (though often overlapping) concepts [27, 11, 4, 33, 23, 36, 24]. While there is undoubtedly heterogeneity in these different usages, there are also recurring themes—such as the relational, contextual, and contested nature of meaning. In this paper, we are primarily concerned with what we call “human meaning,” to distinguish our usage from debates about whether LLMs really “understand” the semantic meaning of the text they generate [6, 50]. We take human meaning to have two components: cultural and experiential.
Cultural meaning. This is the endorsement of certain symbols as significant within the context of a given community. These symbols could include cultural emblems (such as the Union Jack or the Olympic rings), behaviors (such as bowing or raising one’s middle finger), rituals (such as a marriage ceremony or graduation procession), events (such as the Battle of Waterloo or the Moon landing), or works of art (such as James Joyce’s Ulysses or Picasso’s Guernica). This is how the term is typically used in humanities and social science fields such as literary studies [32], history [17], and cultural anthropology [27]. This usage is consistent with other usages of “culture” in contemporary AI [21, 113, 86, 1, 44]. For our purposes, this is the kind of meaning we can most directly make legible with existing frameworks from the humanities and social sciences.
@@ 3 Representing Meaning via Thick Description
In this paper, our primary reference point for meaning is Clifford Geertz’s (1973) account of “thick description [27], which provides a framework for the interpretation of symbols and behaviors within their cultural context to uncover their layers of significance. Methodologically, it is a bedrock of qualitative approaches in the social sciences [87, 64, 58]; theoretically, it is a unifying framework for influential schools of thought in disciplines such as literary criticism [32], history [17], political science [108], philosophy [98], and more. It is arguably one of the most influential frameworks in the humanities and social sciences. [^1] According to Google Scholar’s citation count [20 May 2025], Geertz (1973) has ¿100k citations. For comparison, foundational frameworks in related disciplines published around the same time include: Kahneman & Tversky’s (1979) prospect theory in the behavioral sciences (89k citations) and Granovetter’s (1973) weak ties social network analysis in the quantitative social science (76k citations). Thick description offers a conceptual and methodological bridge between cultural and experiential meaning, as well as a guide for how human meaning can be rendered legible—both to human interpreters and computational systems—at scale.
Geertz proposed thick description as a means of clarifying the goal of ethnographic fieldwork in cultural anthropology. The aim is not just to provide a surface-level account of physical behaviors, rituals, stories, or cultural norms (i.e., thin description). Rather, it is to provide an interpretation of the symbolic dimensions of a given behavior or practice—to analyze its role in providing an individual with a larger sense of meaning and purpose in the context of their broader community (i.e., thick description). Geertz contended that this approach to the study of culture is “not an experimental science in search of law but an interpretive one in search of meaning” [27].
@@
By contrast, a canonical instance of thick description is Malinowski’s (1922) investigation of the Kula Ring [57, 29]. The Kula Ring refers to an elaborate ritual in which inhabitants of the Trobriand Islands crossed treacherous swathes of water in small boats to trade necklaces and armbands which in themselves had no obvious utility: the necklaces and armbands were never actually worn. Malinowski’s account showed how this ostensibly economic activity could only be understood by an analysis of its symbolic dimensions. That is, it cannot be accounted for by an analysis of the transactions through an externally observable value metric (such as money). Rather, it requires an analysis of the meaning assigned to such transactions internally by the people participating in them. The Kula necklaces and armbands were symbols of achievement and social standing. Comparably, an observer attempting to understand the value of a Super Bowl trophy in terms of how much water or soup it holds will likely miss important aspects of the broader cultural activity it represents.
As a theory of meaning, thick description operationalizes meaning as a relationship mediating a symbol (including artifacts such as a Super Bowl trophy or the Bible; actions such as attending a protest or making a particular hand gesture; and mental states such as belief in the sanctity of marriage or satisfaction with clearing one’s email inbox) and its sociocultural context. Influenced by Wittgenstein (1953), Geertz claimed that meaning is not an intrinsic property of an individual mind but only exists within the broader context of who is interpreting a given symbol, the background or experience they bring, and the time and place in which the interpretation occurs [27, 110].
@@ 3.1 Thick and Thin Description in Sociotechnical Systems
In each of these cases, there is a more abstract notion of “what really matters” that gets consolidated into a thin signal which is legible at scale. We contend that employing thick description can help build systems that can represent something closer to the meaningful target notion.
@@ 4.1 Meaning Only Exists in Context
The thrust of our proposal is not that qualitative representations are better than quantitative ones. It is that in order to represent meaning, we need to represent context. There are many ways to do this, with qualitative thick description being a canonical one. But for any given implementation strategy, the standard is not whether it is qualitative or quantitative—but how well it retains information about context in which the meaningful experience or action originally occurred.
This context-dependence of meaning holds across a range of domains and definitions of “meaning.” For example, words are meaningful, but only when used in the context of a given language; a swastika is a meaningful symbol, but its meaning changes when found in Western Europe versus a Buddhist country; and marriage ceremonies can be meaningful, but what specific aspects of the ceremony will be perceived as meaningful depends on the culture in which it is taking place—not to mention the individuals getting married. The difficulty is heterogeneity. What kind of context needs to be represented in order to capture meaning is underconstrained. Representing context is not simply a matter of listing proxy variables, like country, language, or ethnicity, nor can it be encapsulated in a feature set of norms, preferences, or taboos [113, 44, 85]. This is why standardization is antithetical to the representation of human meaning.
One means of translating between thinner and thicker representations is via novel methods for qualitative coding, in which subjective interpretations are condensed into a numerical value by a human analyst [87]. Indeed, the question of translating between thinner and thicker descriptions can be reframed in terms of such coding: How do we create domain-general methods for coding qualitative contextual factors into numerical values? Traditionally, qualitative coding schemes are bespoke for a given dataset—a human researcher must develop, test, and train other researchers to use them. However, with LLMs it may be possible both to develop a unique coding scheme for relevant aspects of a given context and then to employ the newly created scheme [49, 88]. This could help avoid the mandate for standardization by allowing for an ad hoc process of evaluation, adaptable to an arbitrarily large number of contexts, which nonetheless yields numerical values.
@@ 4.2 There Is No Single Source Of Truth for Meaning
The point of codifying meaning is not that there are “right” and “wrong” answers; it is that there are no easy answers and that any oversimplification necessarily omits crucial information about human meaning. The same symbol, artifact, or experience can carry different meanings depending on who is interpreting it and in what context. Systems for representing meaning should maintain this plurality rather than collapsing diverse interpretations into singular representations.
For example, consider the case of a student for whom we have comprehensive educational data, including standardized test scores, demographic information, personal essays, psychological evaluations by third-party experts (think: an augmented college admissions application). One evaluator might interpret an application as providing evidence of technical focus and emerging leadership in STEM fields. Another might view it as the story of a creative, emotionally attuned individual with potential in design or writing. A third might focus on signs of perfectionism or risk-aversion. Each interpretation is grounded in the same underlying data, yet each foregrounds different elements of meaning. This highlights a core difficulty: even when the input is constant, thick descriptions can diverge depending on the interpretive lens. The goal, then, is not to eliminate this divergence but to build systems that can recognize and represent such pluralism. How do we apply interpretive judgment at scale in ways that are consistent, fair, and reflexive?
Recent work shows how LLMs can be used to reckon with this kind of ambiguity, specifically to identify the presence of dog whistles [61, 53]. These are a form of coded communication that uses covert social or ideological signals to convey a generally acceptable meaning to a wider audience while conveying a more controversial or offensive meaning to a targeted audience. This leads to conflicting, audience-dependent layers of meaning. This work shows that LLMs are capable of word-sense disambiguation by analyzing contextual usage, distinguishing between benign and coded meanings, giving a preliminary instance of how LLMs can be used to support interpretive pluralism.
@@ 4.3 Both Lived Experience and Critical Distance Are Crucial
Expert judgments from trained humanists and social scientists are useful, but they must be grounded in the lived experience of the individuals they are trying to understand. Geertz (1983) distinguishes between the concepts of experience-near (informal usages that reflect a subjective perspective) and experience-distant (formal usages that reflect an analytical or scientific perspective) [28]. For example, a gambler and an AI researcher might describe the same situation in different terms (e.g., experience-near: slot machine, hitting it big; experience-distant: n-armed bandit, sparse rewards). Both specialized scholarly knowledge (experience-distant) and lived cultural experience (experience-near) are necessary for robust representations of meaning. Neither should be privileged at the expense of the other, and systems should facilitate integration across these different forms of expertise.

### v2

# Meaning Is Not A Metric: Using LLMs to make cultural context legible at scale

@@ Abstract
This position paper argues that large language models (LLMs) can make cultural context, and therefore human meaning, legible at an unprecedented scale in AI-based sociotechnical systems. We argue that such systems have previously been unable to represent human meaning because they rely on thin descriptions (numerical representations that enforce standardization and therefore strip human activity of the cultural context which gives it meaning). By contrast, scholars in the humanities and qualitative social sciences have developed frameworks for representing meaning through thick description (verbal representations that accommodate heterogeneity and retain contextual information needed to represent human meaning). The verbal capabilities of LLMs now provide a means of at least partially automating the generation and processing of thick descriptions, offering new ways to deploy them at scale. We argue that the problem of rendering human meaning legible is not just about selecting better metrics but about developing new representational formats based on thick description. We frame this as a crucial direction for the application of generative AI and identify five key challenges: preserving context, maintaining interpretive pluralism, integrating perspectives based on lived experience and critical distance, distinguishing qualitative content from quantitative magnitude, and acknowledging meaning as dynamic rather than static.
@@ 1 Introduction
The premise of this paper is that it is important, but difficult, to develop technologies that are expressly designed to support meaningful human experience. Our position aims to articulate how AI technologies can be used not just to yield superficial increases in convenience, but to offer something deeper: tools for helping people tap more directly into what really matters. This is not an inevitable outcome of technological progress. Dystopian fiction is full of cases in which a hypothetical society has advanced far beyond our own, only to find that their sophisticated technology has had a detrimental effect on human experience. Of course, the easy part is proposing that technology should have positive societal impacts; the difficult part is defining “what really matters.”
For example, consider the case of social media. Initially these platforms were presented as tools for human social connection [10]. Over time, however, they came to reflect the more superficial dynamics of engagement metrics; content was prioritized not to facilitate more profound social interactions, but on the basis of likes, views, and shares [112, 31, 70]. This led to a wide range of negative societal outcomes, such as rampant misinformation [60, 115], increased feelings of loneliness [85, 110], addictive behavior [41], and other psychological and social costs [23].
In response, technologists, social scientists, and activists have proposed other ways to prioritize content on these platforms which would, in one form another, allow us to more directly identify and amplify what really matters [35, 103, 17, 62]. A representative example is the Time Well Spent movement (now the Center for Humane Technology) led by Tristan Harris [35]. The goal of this movement was to develop technology that supports “meaningful” social connections and “choices we cherish” [79]. In response, at least one major social media platform alleged to have changed their recommendation algorithm to prioritize these kind of “meaningful” interactions [105]. However, it was disputed whether this effort succeeded, or simply substituted one superficial metric for another [79, 70].
@@
We contend that in order to develop sociotechnical systems that can increase time well spent or reduce drudgery, we first need a way of systematically codifying what kind of experiences humans find meaningful. We approach this problem—how to design sociotechnical systems that support human meaning—as one of representation. The core of our argument is that this issue has been difficult to address not simply because we have been using the wrong metrics, but because “metrics” (as we traditionally conceive of them) are the wrong representational format. The formats we typically use for measuring behavior at scale exclude key contextual information necessary for representing human meaning. The primary aim of this paper will be to describe why that is the case and how we can begin to address it.
Our position is that we can use LLMs to make human meaning legible at an unprecedented scale; that “thin” metrics alone will not be able to do this, because they cannot adequately represent cultural context; and that developing “thick” representational formats might eventually allow our society’s most powerful and impactful sociotechnical systems to support the heterogeneous range of experiences that really matter to people. We argue that such systems have previously been unable to represent human meaning because they typically rely on thin description: numerical representations that enforce standardization, and therefore strip human activity of the broader context that gives it meaning [9, 76, 65, 96]. By contrast, scholars in the humanities and qualitative social sciences have developed robust frameworks for representing meaning based on thick descriptions: verbal representations that allow for heterogeneity, and therefore retain crucial contextual information necessary for analyzing cultural and experiential meaning [28, 117, 94, 89, 62]. These methods provide a way of codifying meaning—but they are difficult to deploy at scale.
@@ 2 The Legibility of Meaning
In his classic political science text Seeing like a State, James C. Scott introduces the concept of legibility [96]. This term describes the connection between a sociotechnical system (in Scott’s case, a political state) and its information processing capacities. Political states cannot represent the goals and desires of their constituents directly. Simply asking each person what they want would result in too much heterogeneity of response. Rather, the state must develop standardized metrics that can be applied across the whole population. These help the state understand the overall trends and dynamics of the populace. However, as Scott argues, by enforcing this standardization the state makes some kinds of information legible—while rendering other kinds illegible.
Drawing on Scott’s concept, we frame the main challenge of this paper as how to make human meaning legible at scale. What it means for something to be “legible” to a given system is that it can be perceived, represented, or processed within the system’s information architecture. For example, in a music streaming platform, only certain aspects of listener behavior are legible [97]. The system can “see” how many clicks a user makes before settling on a song or playlist, whether they range between genres or tend to stick with their favorite one, and which artists they listened to most—the kind of information found in Spotify’s yearly “wrapped” summaries. However, the system cannot “see” whether the user is listening to music while doing work versus socializing. It cannot see the physiological or psychological effects music has on the listener, nor whether the listener likes a given song because of the beat, the lyrics, or the artistic virtuosity. What is legible to a system (i.e., what it “sees”) constrains and shapes the kind of inferences it can make about its users. This in turns affects the content that is prioritized, recommended, and amplified on the platform.
To clarify: we are not concerned with how to make human meaning legible within AI models themselves, but rather how AI models can be used to render new kinds of information legible within sociotechnical systems. We use the term sociotechnical system to refer to any large-scale configuration of people and machines that cannot be understood in terms of human or technological dynamics alone but that depend on a complex entanglement of their interactions [42, 5, 109]. This includes many systems that have a major impact on society, such as political states, social media platforms, corporations, and institutions for healthcare, education, and science. As AI technologies become increasingly integrated into these societally impactful systems, how can we use these technologies to make new kinds of information (such as human meaning) legible at scale?
@@ 2.1 What Do We Mean By “Meaning”?
The term “meaning” has been used by many different people, across many disciplines, to refer to many different (though often overlapping) concepts [28, 11, 4, 34, 24, 38, 25, 62]. While there is undoubtedly heterogeneity in these different usages, there are also recurring themes—such as the relational, contextual, and contested nature of meaning. In this paper, we are primarily concerned with what we refer to as “human meaning,” to distinguish our usage from debates about whether LLMs really “understand” the semantic meaning of the text they generate [6, 52]. We take human meaning to have two components: cultural and experiential.
Cultural meaning. This is the endorsement of certain symbols as significant within the context of a given community. These symbols could include cultural emblems (such as the Union Jack or the Olympic rings), behaviors (such as bowing or raising one’s middle finger), rituals (such as a marriage ceremony or graduation procession), events (such as the Battle of Waterloo or the Moon landing), or works of art (such as James Joyce’s Ulysses or Picasso’s Guernica). This is how the term is typically used in humanities and social science fields such as literary studies [33], history [18], and cultural anthropology [28]. This usage is consistent with other usages of “culture” in contemporary AI [54, 22, 122, 93, 1, 46]. For our purposes, this is the kind of meaning we can most directly make legible with existing frameworks from the humanities and social sciences.
@@ 3 Representing Meaning via Thick Description
In this paper, our primary reference point for meaning is Clifford Geertz’s account of “thick description” [28], which provides a framework for the interpretation of symbols and behaviors within their cultural context to uncover their layers of significance. Methodologically, it is a bedrock of qualitative approaches in the social sciences [94, 71, 64]; theoretically, it is a unifying framework for influential schools of thought in disciplines such as literary criticism [33], history [18], political science [117], and philosophy [107]. It is one of the most influential frameworks in the humanities and social sciences. Thick description offers a conceptual and methodological bridge between cultural and experiential meaning, as well as a guide for how human meaning can be rendered legible—both to human interpreters and computational systems—at scale.
Geertz proposed thick description as a means of clarifying the goal of ethnographic fieldwork in cultural anthropology. The aim is not just to provide a literal account of physical behaviors, rituals, stories, or cultural norms (i.e., thin description). Rather, it is to provide an interpretation of the symbolic dimensions of a given behavior or practice—to analyze its role in providing a larger sense of meaning and purpose in the context of a broader community (i.e., thick description). Geertz contended that this approach to the study of culture is “not an experimental science in search of law but an interpretive one in search of meaning” [28].
@@
By contrast, a canonical instance of thick description is Malinowski’s investigation of the Kula Ring [63, 30]. The Kula Ring refers to an elaborate ritual (now largely defunct) in which inhabitants of the Trobriand Islands crossed treacherous swathes of water in small boats to trade necklaces and armbands which in themselves had no obvious utility: the necklaces and armbands were never actually worn. Malinowski’s account showed how this ostensibly economic activity could only be understood by an analysis of its symbolic dimensions. That is, it cannot be accounted for by an analysis of the transactions through an externally observable value metric (such as money). Rather, it requires an analysis of the meaning assigned to such transactions internally by the people participating in them. The Kula necklaces and armbands were symbols of achievement and social standing. Comparably, an observer attempting to understand the value of a Super Bowl trophy in terms of how much water or soup it holds will likely miss important aspects of the broader cultural activity it represents.
As a theory of meaning, thick description operationalizes meaning as a relationship mediating a symbol (including artifacts such as a championship trophy or the Bible; actions such as attending a protest or making a particular hand gesture; and mental states such as belief in the sanctity of marriage or satisfaction with clearing one’s email inbox) and its sociocultural context. Influenced by Wittgenstein [119], Geertz claimed that meaning is not an intrinsic property of an individual mind but only exists within the broader interpretive context: the background or experience the interpreter brings, and the time and place in which the interpretation occurs [28].
@@ 3.1 Thick and Thin Description in Sociotechnical Systems
In each of these cases, there is an abstract notion of “what really matters” that gets consolidated into a thin signal which is legible at scale. Other authors have referred to this as the problem of “full-stack” alignment [62]. We contend that employing thick description can help build systems that can represent something closer to the meaningful target notion.
@@ 4.1 Meaning Only Exists in Context
The thrust of our proposal is not that qualitative representations are better than quantitative ones. It is that in order to represent meaning, we need to represent context. There are many ways to do this, with qualitative thick description being a canonical one. But for any implementation strategy, the standard is not whether it is qualitative or quantitative—but how well it retains information about the context in which the experience or action originally occurred.
This context-dependence of meaning holds across a range of domains and definitions of “meaning.” For example, words are meaningful, but only when used in the context of a given language; a swastika is a meaningful symbol, but its meaning changes when found in Western Europe versus a Buddhist country; and marriage ceremonies can be meaningful, but what aspects of the ceremony are important depends on the culture in which it is taking place—not to mention the individuals getting married. The difficulty is heterogeneity. What kind of context needs to be represented in order to capture meaning is underconstrained. Representing context is not simply a matter of listing proxy variables, like country, language, or ethnicity, nor can it be encapsulated in a feature set of norms, preferences, or taboos [122, 46, 92]. Such standardization is antithetical to the representation of human meaning.
@@ 4.2 There Is No Single Source Of Truth for Meaning
The point of codifying meaning is not that there are “right” and “wrong” answers; it is that there are no easy answers and that any oversimplification necessarily omits crucial information about human meaning. The same symbol, artifact, or experience can carry different meanings depending on who is interpreting it and in what context. Systems for representing meaning should maintain this plurality rather than collapsing diverse interpretations into unitary variables.
Even when the input is constant, thick descriptions can diverge depending on the interpretive lens. The goal, then, is not to eliminate this divergence but to build systems that can recognize and represent such pluralism. How do we apply interpretive judgment at scale in ways that are consistent, fair, and reflexive? Recent work shows how LLMs can be used to reckon with this kind of ambiguity, specifically to identify the presence of dog whistles [67, 58]. These are a form of coded communication that uses covert social or ideological signals to convey a generally acceptable meaning to a wider audience while conveying a more controversial or offensive meaning to a targeted audience. This leads to conflicting, audience-dependent layers of meaning. This work shows that LLMs are capable of word-sense disambiguation by analyzing contextual usage, distinguishing between benign and coded meanings, giving a preliminary instance of how LLMs can be used to support interpretive pluralism.
@@ 4.3 Both Lived Experience and Critical Distance Are Crucial
Expert judgments from trained humanists and social scientists are useful, but they must be grounded in the lived experience of the individuals they are trying to understand. Geertz distinguishes between the concepts of experience-near (informal usages that reflect a subjective perspective) and experience-distant (formal usages that reflect an analytical or scientific perspective) [29]. For example, a gambler and an AI researcher might describe the same situation in different terms (e.g., experience-near: slot machine, hitting it big; experience-distant: n-armed bandit, sparse rewards). Both specialized scholarly knowledge (experience-distant) and lived cultural experience (experience-near) are necessary for robust representations of meaning. Neither should be privileged at the expense of the other. Systems should facilitate integration across these different forms of expertise.